OpenStudy (anonymous):

SOMEONE PLEASE HELP ALGEBRA 2

OpenStudy (anonymous):

Convert the equation you arrived at in question 2 into slope-intercept form. Make sure to include all of your work. For question two i got x + 1y = 0

OpenStudy (anonymous):

should i explain question two first?

OpenStudy (anonymous):

You don't have to, just thinking.

OpenStudy (anonymous):

Slope-intercept is y = mx + b where m is slope and b is y-intercept. Essentially solve for y. x + 1y = 0 x + y = 0 y = 0 - x y = -x + 0 y = -x

OpenStudy (anonymous):

my coordinates were (1,2) & (0,3) I used (0,3) for question two

OpenStudy (anonymous):

the equations you got are incorrect, unfortunately

OpenStudy (anonymous):

(0, 3) is not a point on the line y = -x

OpenStudy (imstuck):

slope intercept is solving for y and having it in y = mx + b. your equation is not correct for the point-slope form.

OpenStudy (anonymous):

neither is (1, 2)

OpenStudy (anonymous):

okk look imma explain question number two

OpenStudy (anonymous):

2. Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. Launch area (1,2) Point A (0,3) For the slope i got 1 Standard Form - Ax + By = C i got x + 1y = 0

OpenStudy (anonymous):

did i get it right?

OpenStudy (anonymous):

so to get from (1, 2) to (0, 3) what does the spacecraft have to do? that is, how far up or down and how far left or right does it have to move?

OpenStudy (anonymous):

go up

OpenStudy (anonymous):

how much?

OpenStudy (anonymous):

1 left 1 up ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

sorry, my pic is wrong... i'll repost with correct

OpenStudy (anonymous):

did i get number two right?

OpenStudy (anonymous):



OpenStudy (anonymous):

but now you have to adjust for which points the line actually goes through. the slope = rise/run = 1/-1 = -1 = m y=-x+b plug in a known point and solve for b (the y-intercept) or realize that one of the points is the y-intercept (hint, when x = 0). then you'll have slope-intercept form. to get standard form, simply move the x and y to one side of the equation and leave any numbers not associated with x or y on the other side.

OpenStudy (anonymous):

rise = 1 because you go up 1 and run is -1 because you go left 1 (left is the negative direction for the x axis).

OpenStudy (anonymous):

so my slope intercept form would be y = -1(1- 2)

OpenStudy (anonymous):

bc i used the point (1,2)

OpenStudy (anonymous):

no... y = mx+b is slope-intercept form but you need numbers for m and for b. what is the value of m?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

good. now do you know the value of b?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

nope. do you know how to find the value of b?

OpenStudy (anonymous):

no please help

OpenStudy (anonymous):

sure! so first we rewrite the slope-intercept form witth what we know (m, at this point). then we get: y = -1*x + b or y = -x +b now we use a point on the line to help us find the value of b. do you know a point on the line? if so, which one would you like to use?

OpenStudy (anonymous):

(1,2)

OpenStudy (anonymous):

okay, so (1, 2). x = 1 and y = 2. we plug those values into or equation and then solve for b. like this: y = -x +b => (2) = -(1) + b => 2 = -1 + b add 1 to both sides to get 3 = b so our slope-intercept form is y = -x + 3 do you follow all of that?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

you can always check that it works for the points you know on the line. for example, when x = 0, what does the equation tell us y should be?

OpenStudy (anonymous):

0 right?

OpenStudy (anonymous):

nope. put in 0 for x in the equation we just got.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

y = 0 + 3

OpenStudy (anonymous):

there you go! now is the point (0, 3) a point on the line?

OpenStudy (anonymous):

yesss!

OpenStudy (anonymous):

great! what about when x = 1... what is y? again, use the equation.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

3 ?

OpenStudy (anonymous):

nope... y = -(1) + 3 = ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

i didnt use the negative thats why

OpenStudy (anonymous):

there you go! now is (1, 2) a point on the line?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

thank you so much!! you are a very big help <3

OpenStudy (anonymous):

so you have the right slpe-intercept form for your line!

OpenStudy (anonymous):

do you know how to find the standard form?

OpenStudy (anonymous):

Yes i do

OpenStudy (anonymous):

so what is the standard form you got?

OpenStudy (anonymous):

i had got x + 1y = 0

OpenStudy (anonymous):

so you can check that one too. pick a point on the line.

OpenStudy (anonymous):

(0,3)

OpenStudy (anonymous):

so we plug in 0 for x and 3 for y and see if we get a true statement: (0) + 1(3) = 0 + 3 = 3 \(\ne\) 0 so you don't have the correct standard form. do you want to know how to get the correct standard form?

OpenStudy (anonymous):

no please help

OpenStudy (anonymous):

Launch area (1,2) point A (0,3) find the slope slope(m) = (y2 - y1) / (x2 - x1) (1,2) x1 = 1 and x2 = 2 (0,3) x2 = 0 and y2 = 3 now lets sub m = (3 - 2) / (0 - 1) m = 1/1 = 1 if we use (1,2) the equation is x = 1 if we use (0,3), the equation is x = 0 standard form : x + 1y = 0

OpenStudy (anonymous):

or would it be x + 1y = 3

OpenStudy (anonymous):

start with the slope-intercept form we found that we know works... y = -x + 3 now we simply get x and y on the same side. to do this, add x to both sides... x + y = x + (-x) + 3 => x + y = 3 and there is the standard form. you can check your points to make sure it works. check (0, 3) => (0) + (3) = 0 + 3 = 3 check! check (1, 2) => (1) + (2) = 1 + 2 = 3 check!

OpenStudy (anonymous):

yes! so i got it right the second time lol

OpenStudy (anonymous):

yep, great! just follow those steps and you'll be good to go!